# Gravity versus astrophysics in black hole images and photon rings: equatorial emissions and spherically symmetric spacetimes

### v1

@@ Abstract
Context. The Event Horizon Telescope collaboration released in 2019 the first horizon-scale images of a black hole accretion flow, opening a novel route for plasma physics comprehension and gravitational tests. Although the present unresolved images deeply depend on the astrophysical properties of the accreted matter, general relativity predicts that they contain highly lensed observables, the so-called ‘photon rings’, embodying the effects of strong-field gravity.
Aims. Focusing on the particular case of the supermassive black hole M87* and adopting a geometrically thin, equatorial disc as a phenomenological configuration for the accreting matter, our goal is to study the degeneracy of spacetime curvature and of physically-motivated emission processes on plane-of-sky EHT-like images observed at 230 and 345 GHz.
Methods. In a parametric framework, we simulate adaptively ray-traced images using the code GYOTO in various spherically-symmetric spacetime geometries, for a comprehensive class of disc velocities and a library of realistic synchrotron emission profiles. We then extract the width and the peak position of 1D intensity cross sections on the direct image and the first photon ring.
Results. We show that, among the investigated quantities, the most appropriate observables to probe the geometry are the peak positions of the first photon ring. Small geometric deviations can be unequivocally detected, regardless of the motion of the disc, ranging from a Keplerian to a radially infalling one, if the black hole mass-to-distance estimate is accurate up to around 2%; the current uncertainty of 11% being sufficient just to access extreme deviations.
Conclusions. The equatorial set-up of this paper, favoured by present EHT observations of M87*, is adapted to model future measurements at higher observing frequencies, where absorption effects are negligible, and with higher resolution, indispensable to resolve the photon rings. Additional work is needed to investigate if our conclusions hold for more realistic disc configurations.
@@ 1 Introduction
The existence of black holes (BHs) is a direct consequence of general relativity (GR), the standard theory of gravity, and their first solution was found by Karl Schwarzshild (Schwarzschild 1916) a few weeks after the publication of the GR foundations. A cardinal generalisation to rotating black holes, published by Roy Kerr (Kerr 1963), represents the only [^1] With the astrophysically pertinent assumption of null electric charge. stationary solution satisfying the hypotheses of the no-hair theorem (Israel 1967; Carter 1971; Hawking 1972; Robinson 1975) and a Kerr black hole is entirely described by its mass and angular momentum.
@@
The definition of a black hole relies on the presence of an immaterial boundary, the event horizon, that can be crossed only in one direction, towards the BH: not even light can escape it. Also, according to the cosmic censorship conjecture (Penrose 1969), for a Kerr or a Schwarzschild black hole, the event horizon hides observationally inaccessible incomplete null geodesics that may be associated with a curvature singularity. On top of that, the interior of the Kerr spacetime contains closed timelike curves that correspond to a causality violation (Carter 1968). These predictions are controversial and they might mark the limit of the applicability domain of general relativity; it is then justified to continue testing this centennial theory, already thoroughly corroborated for weak gravitational fields, in the extreme conditions of the strong-field regime begotten by black holes.
@@
Since a black hole is not a luminous source itself, its presence is inferred from the influence it has on its environment, which has been revealed by recent revolutionary astronomical observations. The first detection of a gravitational wave signal from the merger of two black holes was announced in 2016 by the LIGO-Virgo collaboration (LIGO Scientific Collaboration and Virgo Collaboration 2016). Two other great achievements, both by GRAVITY in 2018, were the astrometric measurements, close to the supermassive black hole (SMBH) Sagittarius A*, of the orbital motion of the S stars (GRAVITY collaboration 2018a) and of the high states, or flares, of variable near-infrared emission (GRAVITY collaboration 2018b). Finally, the first breakthrough images of black hole accretion flows, probing the strong-field gravity region just above the event horizon, has been released by the Event Horizon Telescope (EHT) in 2019 for M87* (The Event Horizon Telescope Collaboration 2019a).
@@
These images, although compatible with the GR-predicted observational appearance of a Kerr black hole, do not exclude alternative spacetimes (The Event Horizon Telescope Collaboration 2019d; Vincent et al. 2021; Gralla 2021). However, consistency tests of the Kerr-hypothesis will become more stringent with the upgrades of the Next Generation EHT (ngEHT, Johnson et al. 2023; Doeleman et al. 2023; Ayzenberg et al. 2023) and, particularly, with future space-based very-long-baseline interferometers like the proposed mission Black Hole Explorer (BHEX, Johnson et al. 2024; Hudson et al. 2023), that will achieve sufficient angular resolution to access the most gravitationally affected features of the image, namely the photon rings.
@@
It has already been widely shown that the visual aspect of the accretion flow depends both on the spacetime geometry and the properties of the emission, and that the observational appearance of M87*, under the limited instrumental resolution of the EHT, can be mimicked by non-Kerr objects (for instance, Vincent et al. 2021). In this paper we do not attempt to model the limitations of any particular existing or future observing instrument, assuming that a perfect image-domain reconstruction can be obtained. Thus, we focus on the degeneracy between gravity and astrophysics in the images that is fundamental and theoretical in nature. Hence, it is crucial to identify observable components of the images which allow to disentangle these effects. In this work we focus on photon rings, that is, lensed secondary images of the accretion disc superposed on top of the primary image and formed by photons executing at least half-loop around the black hole. In contrast to the primary image and the central depression in brightness, known as the ‘observable shadow’, which are strongly dependent on the astrophysical configuration (The Event Horizon Telescope Collaboration 2019c; Bauer et al. 2022; Kocherlakota & Rezzolla 2022), photon rings are predominantly impacted by the spacetime geometry (Desire et al. 2025).
@@
Photon rings are labelled by the number $n$ of half turns of the photons around the central body. It has been shown that they follow a specific functional form for the Kerr metric not only at high-$n$ (Gralla et al. 2020), but also at $n=1$ (Cárdenas-Avendaño & Lupsasca 2023). Besides, as shown in the latter paper for $n=1$ and elsewhere for $n\geq 2$ (Paugnat et al. 2022), the corresponding interferometric signature is robust under change of the astrophysical configuration for a Kerr black hole. A limitation present in these works, that we would like to overcome here, is the adoption of a three-parameter subset of Johnson’s Standard-Unbounded distribution (Gralla et al. 2020) without physical parameters describing the emission process.
Properties of the photon rings have been used not only to carry out the consistency tests mentioned in the previous paragraph, but also to investigate discriminatory tests of GR, at least for specific classes of metrics and heuristic emission processes:
  • Bauer et al. (2022) and Kocherlakota & Rezzolla (2022) both investigated the properties of the intensity collected on EHT-like images of spherically-symmetric static compact objects surrounded by spherically-infalling emitting matter. Thin-disc models were also studied, as in Eichhorn et al. (2023);
@@
  • Several authors suggested to use Lyapunov exponents to differentiate between spacetime metrics of compact objects that admit separable geodesic equations (Wielgus 2021; Staelens et al. 2023; da Silva et al. 2023; Kocherlakota et al. 2024b). However, since that theoretical feature of photon rings is not directly detectable, one should invoke, for instance, their link with potentially measurable brightness autocorrelations (Hadar et al. 2021; Chesler et al. 2021).
@@
The purpose of this paper is to propose a discriminatory test of general relativity focusing on a parametrised-spherically-symmetric spacetime model of M87*, as in Bauer et al. (2022); Kocherlakota & Rezzolla (2022), and to address the image degeneracy between the nature of the compact object and a large library of physically motivated synchrotron emission profiles for an equatorial disc. We test Keplerian, radially infalling and mixed velocity fields, with a special interest in the first secondary image of the flow, the $n=1$ photon ring. The main restrictions of this preliminary work are the adoptions of a non-rotating black hole and of a geometrically thin accretion disc. Nonetheless, the latter hypothesis may be justified very close to the event horizon both in the framework of classic thick disc accretion models in the plunging region, below the cusp (Abramowicz et al. 1978) and through more recent arguments based on general relativistic magnetohydrodynamic (GRMHD) simulations (Chael et al. 2021). Also, we only simulate instantaneous time-averaged images and we do not explore the associated interferometric visibility as measured by the EHT. These constrictive assumptions can be relaxed in subsequent work by making use of already existing examples for an axisymmetric parametrised spacetime (Konoplya et al. 2016), models for thick accretion flows (Vincent et al. 2022) and astrophysical fluctuations (Lee & Gammie 2021).
This paper is organized as follows. First, Section 2 reviews the definitions of the image components. Then, Sections 3 and 4 detail our models for the compact object and the accretion disc. Next, we describe our simulated images in Section 5 and we analyse some of their distinctive features in Section 6. Finally, we comment on the gravity-astrophysics degeneracy in Section 7 and offer concluding discussion in Section 8.
@@ 2 Schwarzschild image features and observables
Figure 1: Critical curve (dark blue), $n=1$ lensing band (light blue) and inner shadow (brown) of a Schwarzschild black hole on the plane of sky (upper panels) seen at different inclinations and their impact parameters as a function of the polar angle on the screen (lower panels).
In this paper, we consider static and spherically symmetric parametrised black hole metrics, that is, deformations of the vacuum Schwarzschild solution of general relativity. This section reviews the properties of null geodesics in this standard spacetime. However, the notions explored here can be and have been analysed in the spinning case, see Lupsasca et al. (2024) for a general introductory review on Kerr black hole imaging. Here, all the expressions are written in geometrised units ($G=c=1$), so $M$, the mass of the black hole, has the dimension of a length. We recall that the radial position of a Schwarzschild event horizon, in Schwarzschild-Droste coordinates $(t,r,\theta,\varphi)$, is $r_{\mathrm{H}}=2M$.
@@ 2.1 Null geodesics and photon shell
While the equations of geodesic motion in the Schwarzschild [^2] Their study in the Kerr metric was initiated by Carter (Carter 1968) and it was recently revisited and completed (Gralla & Lupsasca 2020b). spacetime were first derived by Schwarzschild (Schwarzschild 1916) and Droste (Droste 1917), the explicit examination of null geodesics governing the trajectory of massless particles like photons is, as far as we know, due to Flamm (Flamm 1916).
The four null geodesic equations can be obtained by means of the expressions translating the symmetries of the spacetime. The two Killing vectors $\boldsymbol{\partial}_{t}$ and $\boldsymbol{\partial}_{\varphi}$ associated with the stationarity and axisymmetry of the Schwarzschild spacetime result in the conserved energy $E$ and conserved angular momentum $L$:
@@
where $p^{\mu}=\mathrm{d}x^{\mu}/\mathrm{d}s$ are the components of the 4-momentum of the photon in Schwarzschild-Droste coordinates, with an affine parameter $s$ that increases along the geodesic towards the future. The energy-rescaled angular momentum $\lambda$, or equivalently the impact parameter $b\coloneqq|\lambda|$, is properly defined for null geodesics reaching an observer at asymptotic infinity, the ones of interest here, for which $E>0$, and it entirely characterises them (Wald 1984). Also, preserving the spherical symmetry confines the photon on a timelike-planar hypersurface and its mass $m=0$ represents the fourth integral of motion.
Excluding worldlines which terminate in the black hole region, there are no stable bound orbits for photons, since $r(s)$ is either a monotonic function or has a single radial turning point. Besides, photons orbit at fixed Schwarzschild-Droste radius $\tilde{r}$ if and only if, in Schwarzschild-Droste coordinates (Hilbert 1917)
@@
As mentioned above, these circular orbits are unstable: at the slightest perturbation, the photon initially at $\tilde{r}$ either falls inside the event horizon or it escapes to infinity. The region of the spacetime spanned by all the circular photon orbits is called the photon sphere [^3] For Kerr BHs we talk of spherical orbits and photon shell (Teo 2003).. A photon winds more and more times, clockwise or anticlockwise according to the sign of $\lambda$, as it approaches the photon sphere, where geodesics loop indefinitely.
@@ 2.2 Critical curve and lensing bands
The critical curve is the theoretical image-plane closed curve depicted by the impact points of null trajectories possessing the same constant of motion $\tilde{\lambda}$ as the bound photon orbit that they asymptotically approach. Simplistically, the critical curve, also referred to as ‘apparent boundary’ (Bardeen et al. 1973), is the projection of the photon sphere on the plane of sky and it delimits two regions according to the impact parameter $b$ of the geodesics. Reversing time’s arrow, one can imagine that the photon is launched from the screen of the observer, at $r_{\mathrm{em}}>\tilde{r}$, towards the black hole: inside the critical curve, that is, in the so-called ‘black hole shadow’ defined by $b<\tilde{b}$, null geodesics cross the horizon, while outside of it, when $b>\tilde{b}$, photon trajectories are not trapped. We stress that ‘black hole shadow’ is an unfortunate naming convention because it does not generally match with the effectively darker region on the image, as it is underlined in Paragraph 2.4. For a Schwarzschild black hole, the critical curve is a circle of radius equal to the critical impact parameter $\tilde{b}=3\sqrt{3}M\simeq 5.196152M$ [^4] In the Kerr case, its angular coordinates on the screen of a distant observer have been first derived by Cunningham and Bardeen (Cunningham & Bardeen 1973) and re-examined in present-day papers (Teo 2003; Johnson et al. 2020). Its cardioid shape, of similar size as the Schwarzschild one, depends on spin and inclination (Falcke et al. 1999).; we represent the critical curve as a dark blue line in Figure 1.
@@
Lensing bands are regions on the observer’s screen that comprise all the impact points reached by photons with constants of motions close to those of a bound trajectory, that is, with $\lambda\approx\tilde{\lambda}$, and that follow nearly bound null orbits (Gralla et al. 2019): these geodesics pass close to the photon shell, without being trapped, in such a way that they can perform several turns around the black hole before leaving to infinity. More precisely, given an equatorial plane passing through the black hole, we say that a light ray belongs to the $n$-th lensing band, with $n\geq 1$, if and only if it intersects the given equatorial plane exactly $n+1$ times on its way to the observer (Chael et al. 2021; Paugnat et al. 2022). If the plane is parallel to the screen, this is equivalent to the statement that the photons of $n$-th lensing band execute $n$ half turns around the black hole before reaching the observer. See Appendix A for an alternative definition of lensing bands.
@@
Lensing bands are perfectly circular for a black hole seen face-on, whereas they are highly deformed for increasing inclinations $i$ of the distant observer [^5] In a Kerr spacetime, they also weakly depend on the spin (Gralla & Lupsasca 2020a)., as shown by the light blue bands of Figure 1. In particular, lensing bands of an inclined screen possess a reflection symmetry [^6] This is only true for spherically symmetric spacetimes (Gralla & Lupsasca 2020a)., they correspond to those of the non-inclined screen along the line perpendicular to the axis of reflection and their width varies according to the polar angle on the screen; see Appendix A for a proof of these statements. Also, the inner edge of the $n$-th lensing band is traced by means of the impact parameters of light rays of order $n$ that asymptotically approach the event horizon, $r\to r_{\mathrm{H}}$, while its outer edge corresponds to the image of a circle of radius $r\to\infty$ on the plane transported by light rays of order $n$, see Figure A.3 in Appendix A. We highlight that in our plots, the outer edge of the lensing band is prescribed by large, but not infinite, values of $r$ according to numerical stop conditions on the disc extent.
@@
The critical curve, as well as the lensing bands defined with respect to a given equatorial plane [^7] See Vincent et al. (2022) and Kocherlakota et al. (2024a) for a discussion on image orders when the crossed shape possesses some thickness. depend exclusively on the geometry of spacetime and are insensitive to the accretion disc properties. Nonetheless, because of the asymptotic character of their precise definitions, they are not the observables that could be directly measured from images of black holes.

### v2

# Gravity versus astrophysics in black hole images and photon rings: Equatorial emissions and spherically symmetric space-times

@@ Abstract
Context. The Event Horizon Telescope (EHT) collaboration released in 2019 the first horizon-scale images of a black hole accretion flow, opening a novel route for plasma physics comprehension and gravitational tests. Although the present unresolved images deeply depend on the astrophysical properties of the accreted matter, general relativity predicts that they contain highly lensed observables, the so-called photon rings, embodying the effects of strong-field gravity.
Aims. Focusing on the particular case of the supermassive black hole M87* and adopting a geometrically thin equatorial disc as a phenomenological configuration for the accreting matter, our goal is to study the degeneracy of space-time curvature and of physically motivated emission processes on plane-of-sky EHT-like images observed at 230 and 345 GHz.
Methods. In a parametric framework, we simulated adaptively ray-traced images using the code GYOTO in various spherically symmetric space-time geometries for a comprehensive class of disc velocities and a library of realistic synchrotron emission profiles. We then extracted the width and the peak position of 1D intensity cross sections on the direct image and the first photon ring.
Results. We show that among the investigated quantities, the most appropriate observables to probe the geometry are the peak positions of the first photon ring. Small geometric deviations can be unequivocally detected regardless of the motion of the disc, ranging from Keplerian rotation to radial infall, if the black hole mass-to-distance estimate is accurate up to around 2%, with the current uncertainty of 11% being just sufficient to access extreme deviations.
Conclusions. The equatorial set-up of this paper, which is favoured by present EHT observations of M87*, is adapted to modelling future measurements at higher observing frequencies, where absorption effects are negligible, and with higher resolution, indispensable to resolving the photon rings. Additional work is needed to investigate if our conclusions hold for more realistic disc configurations.
@@ 1 Introduction
The existence of black holes is a direct consequence of general relativity, the standard theory of gravity, and their first solution was found by Karl Schwarzschild (Schwarzschild 1916) a few weeks after the publication of the general relativity foundations. A cardinal generalisation to rotating black holes, published by Roy Kerr (Kerr 1963), represents the only [^1] This holds in the astrophysically relevant case of null electric charge. stationary solution satisfying the hypotheses of the no-hair theorem (Israel 1967; Carter 1971; Hawking 1972; Robinson 1975), and a Kerr black hole is entirely described by its mass and angular momentum.
@@
The definition of a black hole relies on the presence of an immaterial boundary, the event horizon, that can be crossed only in one direction, towards the black hole. Notably, not even light can escape it. According to the cosmic censorship conjecture (Penrose 1969), for a Kerr or a Schwarzschild black hole, the event horizon also hides observationally inaccessible incomplete null geodesics that may be associated with a curvature singularity. On top of that, the interior of the Kerr space-time contains closed time-like curves that correspond to a causality violation (Carter 1968). These predictions are controversial, and they might mark the limit of the applicability domain of general relativity. It is then justified to continue testing this centennial theory, which has already been thoroughly corroborated for weak gravitational fields, in the extreme conditions of the strong-field regime begotten by black holes.
@@
Since a black hole is not a luminous source itself, its presence is inferred from the influence it has on its environment, which has been revealed by recent revolutionary astronomical observations. The first detection of a gravitational wave signal from the merger of two black holes was announced in 2016 by the LIGO-Virgo collaboration (LIGO Scientific Collaboration and Virgo Collaboration 2016). Two other great achievements, both by GRAVITY in 2018, were the astrometric measurements, close to the supermassive black hole Sagittarius A*, of the orbital motion of the S stars (GRAVITY collaboration 2018b) and of the high states, or flares, of variable near-infrared emission (GRAVITY collaboration 2018a). Finally, the first breakthrough images of black hole accretion flows probing the strong-field gravity region just above the event horizon were released by the Event Horizon Telescope (EHT) in 2019 for M87* (The Event Horizon Telescope Collaboration 2019a).
@@
These images, although compatible with the GR-predicted observational appearance of a Kerr black hole, do not exclude alternative space-times (The Event Horizon Telescope Collaboration 2019d; Vincent et al. 2021; Gralla 2021). However, consistency tests of the Kerr-hypothesis will become more stringent with the upgrades of the Next Generation EHT (ngEHT, Johnson et al. 2023; Doeleman et al. 2023; Ayzenberg et al. 2023) and, particularly, with future space-based very-long-baseline interferometers such as the proposed mission Black Hole Explorer (BHEX, Johnson et al. 2024; Hudson et al. 2023), which will achieve sufficient angular resolution to access the most gravitationally affected features of the image, namely the photon rings.
@@
It has already been widely shown that the visual aspect of the accretion flow depends both on the space-time geometry and the properties of the emission and that the observational appearance of M87*, under the limited instrumental resolution of the EHT, can be mimicked by non-Kerr objects (for instance, Vincent et al. 2021). In this paper we do not attempt to model the limitations of any particular existing or future observing instrument, assuming that a perfect image-domain reconstruction can be obtained. Thus, we focus on the degeneracy between gravity and astrophysics in the images that is fundamental and theoretical in nature. Hence, it is crucial to identify observable components of the images that allow these effects to be disentangled. In this work we focus on photon rings, that is, lensed secondary images of the accretion disc superposed on top of the primary image and formed by photons executing at least a half loop around the black hole. In contrast to the primary image and the central depression in brightness, known as the ‘observable shadow’, which are strongly dependent on the astrophysical configuration (The Event Horizon Telescope Collaboration 2019c; Bauer et al. 2022; Kocherlakota & Rezzolla 2022), photon rings are predominantly impacted by the space-time geometry (Desire et al. 2025).
@@
Photon rings are labelled by the number of half turns, $n$, of the photons around the central body. It has been shown that they follow a specific functional form for the Kerr metric not only at high $n$ (Gralla et al. 2020) but also at $n=1$ (Cárdenas-Avendaño & Lupsasca 2023). Moreover, as shown in the latter paper for $n=1$ and elsewhere for $n\geq 2$ (Paugnat et al. 2022), the corresponding interferometric signature is robust under change of the astrophysical configuration for a Kerr black hole. A limitation present in these works that we would like to overcome here is the adoption of a three-parameter subset of Johnson’s Standard-Unbounded distribution (Gralla et al. 2020) without physical parameters describing the emission process.
@@
Properties of the photon rings have been used not only to carry out the consistency tests mentioned in the previous paragraph but also to investigate discriminatory tests of general relativity, at least for specific classes of metrics and heuristic emission processes. For example, Bauer et al. (2022) and Kocherlakota & Rezzolla (2022) have both investigated the properties of the intensity collected on EHT-like images of spherically symmetric static compact objects surrounded by spherically infalling emitting matter. In this framework, thin disc models have also been studied, as in Eichhorn et al. (2023). Several other authors have suggested the use of Lyapunov exponents to differentiate between the space-time metrics of compact objects that admit separable geodesic equations (Wielgus 2021; Staelens et al. 2023; da Silva et al. 2023; Kocherlakota et al. 2024b). However, since that theoretical feature of photon rings is not directly detectable, one should invoke, for instance, their link with potentially measurable brightness autocorrelations (Hadar et al. 2021; Chesler et al. 2021).
@@
The purpose of this paper is to propose a discriminatory test of general relativity focusing on a parametrised spherically symmetric space-time model of M87*, as in Bauer et al. (2022); Kocherlakota & Rezzolla (2022), and to address the image degeneracy between the nature of the compact object and a large library of physically motivated synchrotron emission profiles for an equatorial disc. We tested Keplerian, radially infalling, and mixed velocity fields, with a special interest in the first secondary image of the flow, the $n=1$ photon ring. The main restrictions of this preliminary work are the adoptions of a non-rotating black hole and of a geometrically thin accretion disc. Nonetheless, the latter hypothesis may be justified very close to the event horizon both in the framework of classic thick disc accretion models in the plunging region, below the cusp (Abramowicz et al. 1978), as well as through more recent arguments based on general relativistic magnetohydrodynamic (GRMHD) simulations (Chael et al. 2021). Also, we only simulated instantaneous time-averaged images, and we did not explore the associated interferometric visibility as measured by the EHT. These constrictive assumptions can be relaxed in subsequent work by making use of already existing examples for an axisymmetric parametrised space-time (Konoplya et al. 2016), models for thick accretion flows (Vincent et al. 2022), and astrophysical fluctuations (Lee & Gammie 2021).
This paper is organised as follows. First, Section 2 provides a review of the definitions of the image components. Then, in Sections 3 and 4 we detail our models for the compact object and the accretion disc. Next, we describe our simulated images in Section 5, and we analyse some of their distinctive features in Section 6. Finally, we comment on the gravity-astrophysics degeneracy in Section 7 and offer a concluding discussion in Section 8.
@@ 2 Schwarzschild image features and observables
Figure 1: Critical curve (dark blue), $n=1$ lensing band (light blue), and inner shadow (brown) of a Schwarzschild black hole on the plane of sky (upper panels) seen at different inclinations and their impact parameters as a function of the polar angle on the screen (lower panels).
In this paper, we consider static and spherically symmetric parametrised black hole metrics, that is, deformations of the vacuum Schwarzschild solution of general relativity. This section reviews the properties of null geodesics in this standard space-time. However, the notions explored here can be and have been analysed in the spinning case (see Lupsasca et al. (2024) for a general introductory review on Kerr black hole imaging). Here, all the expressions are written in geometrised units ($G=c=1$), so $M$, the mass of the black hole, has the dimension of a length. We recall that the radial position of a Schwarzschild event horizon, in Schwarzschild-Droste coordinates $(t,r,\theta,\varphi)$, is $r_{\mathrm{H}}=2M$.
@@ 2.1 Null geodesics and photon shell
While the equations of geodesic motion in the Schwarzschild [^2] Their study in the Kerr metric was initiated by Carter (Carter 1968) and it has recently been revisited and completed (Gralla & Lupsasca 2020b). space-time were first derived by Schwarzschild (Schwarzschild 1916) and Droste (Droste 1917), the explicit examination of null geodesics governing the trajectory of massless particles like photons is, as far as we know, due to Flamm (Flamm 1916).
The four null geodesic equations can be obtained by means of the expressions translating the symmetries of the space-time. The two Killing vectors, $\boldsymbol{\partial}_{t}$ and $\boldsymbol{\partial}_{\varphi}$, associated with the stationarity and axisymmetry of the Schwarzschild space-time result in conserved energy, $E$, and conserved angular momentum, $L$:
@@
where $p^{\mu}=\mathrm{d}x^{\mu}/\mathrm{d}s$ are the components of the 4-momentum of the photon in Schwarzschild-Droste coordinates, with an affine parameter, $s$, that increases along the geodesic towards the future. The energy-rescaled angular momentum, $\lambda$, or equivalently the impact parameter, $b\coloneqq|\lambda|$, is properly defined for null geodesics reaching an observer at asymptotic infinity, the ones of interest here, for which $E>0$, and it entirely characterises them (Wald 1984). Also, preserving the spherical symmetry confines the photon on a time-like-planar hypersurface and its mass, $m=0$, represents the fourth integral of motion.
Excluding worldlines that terminate in the black hole region, there are no stable bound orbits for photons, since $r(s)$ is either a monotonic function or has a single radial turning point. Besides, photons orbit at fixed Schwarzschild-Droste radius, $\tilde{r}$, if and only if, in Schwarzschild-Droste coordinates (Hilbert 1917)
@@
As mentioned above, these circular orbits are unstable: at the slightest perturbation, the photon initially at $\tilde{r}$ either falls inside the event horizon or it escapes to infinity. The region of the space-time spanned by all the circular photon orbits is called the photon sphere. [^3] For Kerr black holes we talk of spherical orbits and photon shell (Teo 2003). A photon winds more and more times, clockwise or anticlockwise according to the sign of $\lambda$, as it approaches the photon sphere, where geodesics loop indefinitely.
@@ 2.2 Critical curve and lensing bands
The critical curve is the theoretical image-plane closed curve depicted by the impact points of null trajectories possessing the same constant of motion, $\tilde{\lambda}$, as the bound photon orbit that they asymptotically approach. Simplistically, the critical curve, also referred to as ‘apparent boundary’ (Bardeen et al. 1973), is the projection of the photon sphere on the plane of sky and it delimits two regions according to the impact parameter, $b$, of the geodesics. Reversing time’s arrow, one can imagine that the photon is launched from the screen of the observer, at $r_{\mathrm{em}}>\tilde{r}$, towards the black hole: inside the critical curve, that is, in the so-called ‘black hole shadow’ defined by $b<\tilde{b}$, null geodesics cross the horizon, while outside of it, when $b>\tilde{b}$, photon trajectories are not trapped. We stress that ‘black hole shadow’ is an unfortunate naming convention because it does not generally match with the effectively darker region on the image, as it is underlined in Paragraph 2.4. For a Schwarzschild black hole, the critical curve is a circle of radius equal to the critical impact parameter, $\tilde{b}=3\sqrt{3}M\simeq 5.196152M$; [^4] In the Kerr case, its angular coordinates on the screen of a distant observer were first derived by Cunningham and Bardeen (Cunningham & Bardeen 1973) and have been re-examined in present-day papers (Teo 2003; Johnson et al. 2020). Its cardioid shape, of similar size as the Schwarzschild one, depends on spin and inclination (Falcke et al. 1999). we represent the critical curve as a dark blue line in Figure 1.
@@
Lensing bands are regions on the observer’s screen that comprise all the impact points reached by photons with constants of motions close to those of a bound trajectory, that is, with $\lambda\approx\tilde{\lambda}$, and that follow nearly bound null orbits (Gralla et al. 2019): these geodesics pass close to the photon shell, without being trapped in such a way that they can perform several turns around the black hole before leaving to infinity. More precisely, given an equatorial plane passing through the black hole, we say that a light ray belongs to the $n$-th lensing band, with $n\geq 1$, if and only if it intersects the given equatorial plane exactly $n+1$ times on its way to the observer (Chael et al. 2021; Paugnat et al. 2022). If the plane is parallel to the screen, this is equivalent to the statement that the photons of $n$-th lensing band execute $n$ half turns around the black hole before reaching the observer. An alternative definition of lensing bands is presented in Appendix A.
@@
Lensing bands are perfectly circular for a black hole seen face-on, whereas they are highly deformed for increasing inclinations, $i$, of the distant observer, [^5] In a Kerr space-time, they also weakly depend on the spin (Gralla & Lupsasca 2020a). as shown by the light blue bands of Figure 1. In particular, lensing bands of an inclined screen possess a reflection symmetry, [^6] This is only true for spherically symmetric space-times (Gralla & Lupsasca 2020a). they correspond to those of the non-inclined screen along the line perpendicular to the axis of reflection and their width varies according to the polar angle on the screen (see Appendix A for a proof of these statements). Also, the inner edge of the $n$-th lensing band is traced by means of the impact parameters of light rays of order $n$ that asymptotically approach the event horizon, $r\to r_{\mathrm{H}}$, while its outer edge corresponds to the image of a circle of radius $r\to\infty$ on the plane transported by light rays of order $n$ (see Figure A.3 in Appendix A). We highlight that in our plots, the outer edge of the lensing band is prescribed by large, but not infinite, values of $r$ according to numerical stop conditions on the disc extent.
@@
The critical curve, as well as the lensing bands defined with respect to a given equatorial plane [^7] Vincent et al. (2022) and Kocherlakota et al. (2024a) offer a discussion on image orders when the crossed shape possesses some thickness. depend exclusively on the geometry of space-time and are insensitive to the accretion disc properties. Nonetheless, because of the asymptotic character of their precise definitions, they are not the observables that could be directly measured from images of black holes.